How many atoms of 12C are there in a 3.50 grams of a pure sample of this particular isotope?
Answer #1

Given mass of 12C = 3.50 g

=> Moles of 12C = mass / atomic mass of 12C = 3.50 g / 12.0 g/mol = 0.2917 mol

1 mole contains 6.022*1023 12C atoms.

=> Number of atomic present in 0.2917 mol 12C = 0.2917 mol * (6.022*1023 12C atoms / 1 mol) = 1.76*1023 atoms (Answer)